Youve collected the following information about Odyssey, Ind. Sales Net income Dividends Total debt Total equity $165,000 $ 12,600 $ 8,300 $66,000 $55,000 What is the sustainable growth rate for the company? (Do not round intermediate calculations. Round your answer to 2 decimal places. ( e.g., 32.16)) Sustainable growth rate If it does grow at this rate, how much new borrowing will take place in the coming year, assuming a constant debt-equity ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16) Additional borrowing What growth rate could be supported with no outside financing at all? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Internal growth rate
Answer #1

a). b = 1 - DPR

= 1 - [$8,300 / $12,600] = 1 - 0.6587 = 0.3413

ROE = Net Income / Total Equity = $12,600 / $55,000 = 0.2291, or 22.91%

Sustainable growth rate = (ROE × b) / [1 - (ROE × b)]

= (0.2291 x 0.3413) / [1 - (0.2291 x 0.3413)]

= 0.0782 / 0.9218 = 0.0848, or 8.48%

b). If the company grows at the sustainable growth rate, the new level of total assets is:

New TA = 1.0848[$66,000 + $55,000] = $131,262.33

To find the new level of debt in the company’s balance sheet, we take the percentage of debt in the capital structure times the new level of total assets. The additional borrowing will be the new level of debt minus the current level of debt. So:

New TD = [D / (D + E)](TA)

= [$66,000 / ($66,000 + $55,000)]($131,262.33) = 0.5455 x $131,262.33 = $71,597.63

Additional borrowing = New TD - Old TD

= $71,597.63 - $66,000 = $5,597.63

c). The growth rate that can be supported with no outside financing is the internal growth rate. To calculate the internal growth rate, we first need the ROA, which is:

ROA = Net Income / Total Assets

= $12,600 / ($66,000 + $55,000) = 0.1041, or 10.41%

Internal growth rate = (ROA × b) / [1 - (ROA × b)]

= (0.1041 x 0.3413) / [1 - (0.1041 x 0.3413)]

= 0.0355 / 0.9645 = 0.0368, or 3.68%